# From Decision to Action in Surgical Autonomy: Multi-Modal Large Language Models for Robot-Assisted Blood Suction

### v1

@@ IV-B User Study on Closeness to Human-Like Behaviour
To assess the similarity to human-like behavior exhibited by the RR, NR, and LRWOC modules during blood suction, a survey questionnaire was conducted involving ten participants, mainly graduate students and senior researchers. A total of thirty-six videos were collected for the survey, divided among three reasoning modules (RR, NR, LRWOC), with each module contributing twelve videos from four environments (three videos per environment), showcasing the blood suctioning task. Participants were presented with pairs of videos and asked, ”If you were the human operator, which of the two videos shown below would you choose to suction the blood pools?”. This forced choice questioning process was repeated to compare all combinations of LRWOC versus NR versus RR.
@@
Although results show a promising stride towards explainable surgical autonomy, a thorough evaluation process including clinical trials is essential to establish the efficacy and safety of LLM-powered robotic systems in real-world surgical applications. Additionally, the development of new training protocols for surgical teams on interacting with and overseeing LLM-enabled autonomous systems, along with intuitive interfaces for surgeons to interact with and override the system’s decisions when necessary, will be vital for adoption.
@@
V Conclusion and Future Work
In this study, we proposed a multi-modal LLM integration in robot-assisted surgery for autonomous blood suction and investigated how the addition of a high-level reasoning unit can influence decision-making and performance. Experiments were conducted to analyze LLM reasoning in comparison to random reasoning and no reasoning modules. Active bleeding and blood clots were introduced to influence decision-making as is also common in highly dynamic surgical settings. Results showed that the presence of a multi-modal LLM as a higher-level reasoning unit can account for these surgical complexities in decision-making and prioritization while also improving the time to suction 50% of the blood to achieve a level of reasoning and explainability previously unattainable in robot-assisted surgeries. The user study showed that incorporating contextual understanding in robotic surgery could bridge the gap between automated procedures and the intuitive decision-making of humans.
While simulation-based environments have certain limitations, they excel in encompassing a comprehensive range of scenarios, including rare but pivotal situations, and are a first critical step to ensuring that the system’s decision-making capabilities are reliable under various conditions. In this study, we assumed that blood pools are separate and independent and the generation speed of OpenAI’s GPT-4V, which led to our system not operating in real-time, served as limitations. The physical experiments demonstrated in our previous work in blood suction [10] show the real-world execution of this task and hence were not the main focus of this study. Future work will investigate mid-sequence task re-planning as a result of interruptions or external disturbances, fine-tuning LLMs using reinforcement learning from human feedback, and exploring the scalability of our approach to other surgical tasks.

### v2

@@ IV-C User Study on Closeness to Human-Like Behaviour
To assess the similarity to human-like behavior exhibited by the RR, NR, and LRWOC modules during blood suction, a survey questionnaire was conducted involving ten participants. The participants, consisting of graduate students and senior researchers with no specialized medical backgrounds, were chosen to assess how well a non-expert-defined context aligns with broader, non-specialist perspectives in evaluating the behavior of the RR, NR, and LRWOC modules. A total of thirty-six videos were collected for the survey, divided among three reasoning modules (RR, NR, LRWOC), with each module contributing twelve videos from four environments (three videos per environment), showcasing the blood suctioning task. Participants were presented with pairs of videos and asked, “If you were the human operator, which of the two videos shown below would you choose to suction the blood pools?”. This forced choice questioning process was repeated to compare all combinations of LRWOC versus NR versus RR.
@@
Although results show a promising stride towards explainable surgical autonomy, a thorough evaluation process including clinical trials is essential to establish the efficacy and safety of LLM-powered robotic systems in real-world surgical applications. Additionally, the development of new training protocols for surgical teams on interacting with and overseeing LLM-enabled systems, along with intuitive interfaces for surgeons to interact with and override the system’s decisions when necessary, will be vital for adoption.
V Limitations and Future Work
While the proposed method demonstrates the effectiveness of multi-modal LLMs in reasoning and decision-making within a simulated environment, several limitations and areas for future work remain, particularly for transitioning to real-world applications in surgical settings.
This study assumed that blood pools are separate and independent, simplifying interactions within the environment. Additionally, the current system operates below real-time performance due to the generation speed of OpenAI’s GPT-4V, which constrains its immediate applicability in time-sensitive clinical tasks. The physical experiments demonstrated in our previous work in blood suction [10] show the real-world execution of this task and hence were not the main focus of this study.
Simulation-based environments, while inherently limited compared to real-world settings, offer the critical advantage of encompassing a wide range of scenarios, including rare but pivotal situations that are difficult to consistently reproduce in physical setups. This simulation-first approach provides a foundation to test and refine the system’s decision-making capabilities across varied conditions, setting the groundwork for application in real surgical environments.
@@ V Limitations and Future Work
Applying this framework to real surgical contexts introduces practical challenges, such as accurate segmentation of blood from live camera images, real-time pose coordination of surgical instruments, and adapting to the dynamic and complex environment of an actual operating room. Future work will address these challenges by refining and validating the framework’s applicability in real surgical tasks, exploring predictive mechanisms and pre-emptive prompts to prioritize time-sensitive actions. To enhance real-time feasibility, model distillation, quantization, and smaller, task-specific models will be investigated, enabling the system to better align with the operational demands of clinical practice.
Ensuring the accuracy and safety of LLM-based decisions is another priority. LLMs are prone to hallucinations, which could impact decision reliability. To mitigate this, future iterations will integrate rule-based checks and domain-specific constraints, as well as feedback loops from medical experts, to improve the system’s robustness and decision accuracy in high-stakes environments. Gathering insights from medical professionals will also help assess the LLM’s alignment with expert-defined surgical priorities, refining its adaptability to clinical needs.
Safety in surgical tasks is essential, and the current framework will be expanded to address this by incorporating force-based thresholds to control tool speed and acceleration, collision detection, and Safe Reinforcement Learning (Safe RL) techniques, such as reward shaping and risk-sensitive policies, to enhance safe tool proximity to sensitive tissues. Additionally, to better handle dynamic scenarios, future work will explore mid-sequence task re-planning to adapt to external disturbances and refine LLM reasoning capabilities through reinforcement learning from human feedback, ultimately improving the framework’s scalability to other surgical and medical tasks.
VI Conclusion
In this study, we proposed a multi-modal LLM integration in robot-assisted surgery for autonomous blood suction and investigated how the addition of a high-level reasoning unit can influence decision-making and performance. Experiments were conducted to analyze LLM reasoning in comparison to random reasoning and no reasoning modules. Active bleeding and blood clots were introduced to influence decision-making as is also common in highly dynamic surgical settings. Results showed that the presence of a multi-modal LLM as a higher-level reasoning unit can account for these surgical complexities in decision-making and prioritization to achieve a level of reasoning and explainability previously unattainable in robot-assisted surgeries. The user study showed that incorporating contextual understanding in robotic surgery could bridge the gap between automated procedures and the intuitive decision-making of humans.